OpenStudy (anonymous):

Simplify (5x)-2 to the 2nd power

OpenStudy (anonymous):

Would the answer be 25x?

OpenStudy (anonymous):

\[(5x)-2 ^2 \]

OpenStudy (anonymous):

or \[(5x)^2\]

OpenStudy (anonymous):

Sorry its supposed to look like the second one except the 2 is negative

OpenStudy (anonymous):

so like this: \((5x)^{-2}\)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[(5x)^{-2}\] To remove the negative exponent, move the \((5x)^{-2}\) to the denominator and make the exponent positive. \(\frac{ 1 }{ (5x)^{2} }\) Expand the exponent (2) to the expression. \(\frac{ 1 }{ 5^{2} x ^{2} }\) Squaring a number is the same as multiplying the number by itself \((5\times5)\). \( \frac{ 1 }{ 5\times5x^2 }\) Multiply 5 by 5 to get 25. \[\frac{ 1 }{ 25x^2 }\]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw :) medal?